OpenStudy (anonymous):

Two boxes are connected to each other. The box on the table weighs 3kg, and on the other end is a suspended 1kg box which is one meter from the floor. The connecting string runs over a pulley. Without taking any friction into account, what is the kinetic energy of the 1kg box just before it hit the floor once the system is released from rest ????

OpenStudy (anonymous):



OpenStudy (anonymous):

force on 3kg box is 9.8N (because 1kg * 9.8m/s)

OpenStudy (anonymous):

the acceleration the 3kg experiences is 9.8N/3kg = 3.26666 m/s^2

OpenStudy (anonymous):

Vf ^2 = Vi^2 + 2ad

OpenStudy (anonymous):

therefore the final velocity on the 3kg box Vf^2 = 0 + 2 (3.26666 m/s^2)(1m)

OpenStudy (anonymous):

Vf = 2.5560 m/s

OpenStudy (anonymous):

the small box experiences the same velocity since their connected by the string

OpenStudy (anonymous):

KE = 1/2 mv^2 KE = 1/2 (1kg)(2.5560m/s)^2

OpenStudy (anonymous):

Hi the answers given to me to choose from are: a) 2,45J b) 4,90J c) 9,80J d) 29,4J e) 39,2J

OpenStudy (anonymous):

it should be a) total PE=KE of 3kg mass + KE of 1kg mass and since they have the same speed 1 x 1 x9.8 = 0.5 x 3 x v^2 + 0.5 x 1 x v^2 since they have the same speed. solving for this you get v^2 = 4.9 KE of the 1kg box is 0.5mv^2=0.5 x 1 x4.9 = 2.95

OpenStudy (tammy18):

According to the principle of conservation of energy, the loss of the gravitational is equal to the kinetic energy gained, thus 1/2 mv^ 2 = mgh, where 1/2 mv^2 is the formula to calculate kinetic energy while mgh is the formula to calculate potential energy. Kinetic energy before hitting the ground = mgh = (1)(10)(1) = 10 N

OpenStudy (souvik):

i am agree with @Matthewszh...used the principle of conservation of energy.. u may solve it like this... and let the acceleration of the 1kg mass is downward...so the 3kg mass has the same acceleration... for 1 kg mass- 1g-T=1*a and for 3 kg mass- T=3*a from this two equation...a=g/4 [g=gravitational acceleration] now let the velocity of the 1 kg mass just before hitting the ground be v so v^2=0+2g/4*1=g/2 so the KE=1/2*1*v^2=g/4=2.45 J